A bank loaned out P120000, part of it at the rate of 8% per year and the rest of the rate of 18% per year.If the interest recieved total P10000,how much was loaned at 8%?

Answer by greenestamps(13200)



A standard formal algebraic solution....

x = amount at 8%
120000-x = amount at 18%

The total interest was 10000:

.08(x)+.18(120000-x)=10000

Solve using basic algebra; I leave that to you.

If a formal algebraic solution is not required, here is a quick way to solve any 2-part mixture problem like this using a bit of mental arithmetic.

(1) Calculate that 10000 interest on a loan of 120000 is a percentage rate of 8 1/3%.
(2) Picture the three percentage rates 8, 8 1/3, and 18 on a number line and observe/calculate that 8 1/3 is (1/3)/10 = 1/30 of the way from 8 to 18.
(3) That means 1/30 of the total was loaned at the higher rate.

ANSWERS: 1/30 of 120000, or 4000 was loaned at 18%; the other 116000 at 8%

CHECK: .08(116000)+.18(4000) = 9280+720=10000